OpenStudy (konradzuse):

Use implicit differentiation to find an equation of the tangent line to the graph at the given point. x + y − 1 = ln(x^10 + y^17), (1, 0)

OpenStudy (konradzuse):

if it's like the other question then it would be y = -x +1 +ln(x^10 + y^17)?

OpenStudy (konradzuse):

then derivation would be d/dx(-x) d/dx(1), ln(x^10 + y^17) and d/dx(x^10 + y^17) because of the chain rule?

OpenStudy (konradzuse):

which should be dy/dx = -1 + \[\frac{10x^9}{x^{10}} + \frac{17x^{16}}{x^{17}}\] +\[10x^9 + 17y^{16}\] ?

OpenStudy (turingtest):

let's just look at the left side first what is the derivative with respect to x of that?

OpenStudy (konradzuse):

I left y alone on the right, should I have no done that?

OpenStudy (konradzuse):

if I didn't it would be 1+ dy/dx

OpenStudy (turingtest):

that's right :) so we have 1+y'= ??? give the derivative of the RHS a try

OpenStudy (konradzuse):

I posted what I thought up above, is that not right?

OpenStudy (turingtest):

well you have it all listed out so it's hard to tell looks like you turned a y into x somehow though....

OpenStudy (konradzuse):

yeah I did oops :P

OpenStudy (turingtest):

\[D_x(\ln u)=\frac{u'}u\]in your case\[u=x^{10}+y^{17}\]so what is \(u'\) ?

OpenStudy (konradzuse):

\[10x^9 + 17y^{16} ?\]

OpenStudy (turingtest):

almost, but you forgot the most important part of implicit differentiation: because y is a function of x, we have to use the chain rule on y, so we get\[10x^9+17y^{16}y'\]that y' (same thing as dy/dx, different notation) is what we are after. Omitting it will give the wrong answer.

OpenStudy (konradzuse):

so it's only for y, and only used for implicit differentiation? What exactly does that term mean?

OpenStudy (konradzuse):

and it just stays as y'? no function?

OpenStudy (turingtest):

y is f(x) in disguise so we could write this y(x) when you take the derivative of some function of y(x) with respect to x that means you have to use the chain rule for example if y(x) is raised to some power the chain rule with respect to x gives\[\frac d{dx}[y(x)]^n=n[y(x)]^{n-1}\cdot y'(x)\]

OpenStudy (turingtest):

\[y'=\frac{dy}{dx}\]they are the same thing, it's just shorthand

OpenStudy (konradzuse):

ic... interesting. So if we have something with respect to something else, no matter what it is we have to use the chain rule.

OpenStudy (konradzuse):

it coudl be y, x, t, u, etc?

OpenStudy (turingtest):

yes, we can always work our way inward toward the dependent variable we are taking the derivative with respect to really this is\[u=y^n\]\[y=\ln t\]then\[\frac{du}{dt}=\frac{du}{dy}\cdot\frac{dy}{dt}=ny^{n-1}\cdot\frac1t=n(\ln t)^{n-1}\cdot\frac1t\]you can always sub back in for y (or whatever intermediate variable you have)at the end

OpenStudy (turingtest):

in the case of your problem, however, subbing in for y in terms of x would be pointless and painful, so we won't do it this time

OpenStudy (konradzuse):

okies :)

OpenStudy (turingtest):

so now we have\[1+y'={10x^9+17y^{16}y'\over x^{10}+y^{17}}\]do you agree?

OpenStudy (konradzuse):

yuh.. At first I guiess I got confused when I split x and y apart, I guess they always stay together as 'u' within ln() and other situations... :p

OpenStudy (turingtest):

exactly, they stay together because they are the "u" in our formula

OpenStudy (turingtest):

so the fancy way to do this is to solve it for y' (which is the slope remember), then plug in the point (1,0) and see what we get however we could also be lazier and just plug in the numbers right now ;) that should make solving for y' a lot easier

OpenStudy (konradzuse):

so it would be 10(1)^9/1^10

OpenStudy (konradzuse):

so 10/1 = 10 = 1 + y' ....... y' = 10-1 = 9

OpenStudy (turingtest):

yep :)

OpenStudy (konradzuse):

didnt' do the other part since it's all 0.

OpenStudy (turingtest):

now we can apply point slope form (hope you remember it) from algebra\[y-y_0=m(x-x_0)\]

OpenStudy (konradzuse):

so now we do the y-y0 = m(x-xo)?

OpenStudy (turingtest):

echoechocecho....

OpenStudy (konradzuse):

ok so that's y-0 = 9(x-1) = y = 9x-9

OpenStudy (turingtest):

and we're done :D

OpenStudy (konradzuse):

:D

OpenStudy (konradzuse):

so when would we solve for y'?

OpenStudy (konradzuse):

also if lets say y was 1 and x was 0 we would have 17(1)^16/(1)^17 * y' right?

OpenStudy (turingtest):

we used y' already, it was our m, remember?

OpenStudy (konradzuse):

oh okay yeah.

OpenStudy (turingtest):

right to your last post

OpenStudy (konradzuse):

you were saying before that was the "fancy way."

OpenStudy (konradzuse):

do the y's just cancel out?

OpenStudy (turingtest):

no, if we \(don't\) plug in the numbers right away, and try to solve for y', that was what I meant by the fancy way, because it is harder and gives a more general result the y' will not cancel out, it should be factorable if you do it right

OpenStudy (konradzuse):

ok because it's 1+y' = (all that stuff) *y' was just curious what happens since there are 2 of them.

OpenStudy (konradzuse):

if we plug in for x and y that doesn't affect the y' on the right , right?

OpenStudy (turingtest):

you would get\[1+y'={10x^9+17y^{16}y'\over x^{10}+y^{17}}\]\[x^{10}+y^{17}+(x^{10}+y^{17})y'=10x^9+17y^{16}y'\]\[x^{10}+y^{17}-10x^9=17y^{16}y'-(x^{10}+y^{17})y'\]

OpenStudy (konradzuse):

but maybe y is always 0 so that explains why the y' cancels out on the right because you're multiplying by 0 but iDk :p

OpenStudy (turingtest):

\[x^{10}+y^{17}-10x^9=[17y^{16}-(x^{10}+y^{17})]y'\]\[y'={x^{10}+y^{17}-10x^9\over17y^{16}-x^{10}-y^{17}}\]notice this gives 9 also, so we're good :)

OpenStudy (turingtest):

the factoring out of the y' can be a little tricky sometimes in these things, so try to focus on that part

OpenStudy (turingtest):

y is not always 0, we cannot treat y as zero until we are ready to solve for y'

OpenStudy (konradzuse):

hmp... so if we just go in and substitute we get rid of the y' on the right? I'm just confused where it goes... I see in this case it's gone because we multiply it by 0, but what if we don't?

OpenStudy (turingtest):

then we could solve it by simple arithmetic say we had 3+y'=y^3y'+x at point (1,2) we could just plug in the numbers now and get 3+y'=8y'+1 7y'=2 y'=2/7 or something like that

OpenStudy (turingtest):

if, somehow, y' did cancel out (like if we had the above evaluated at (1,1) that would tell us that the derivative is undefined there I think... actually I'm not sure, but it doesn't happen often in problems I encounter

OpenStudy (konradzuse):

\[y^{3y'}\] ????

OpenStudy (turingtest):

sorry for my lack of clarity\[3+y'=y^3y'+x\]at (1,2) becomes\[3+y'=(2^3)y'+1\]\[3+y'=8y'+1\]\[7y'=2\]\[y'=2/7\]

OpenStudy (konradzuse):

ah i got it haha. I was konfusing myself with this calculus stuff I forgot about basic math.. I kept thinking it would just be y' all the time iand it would just cancel itself out or be cancelled by 0, thanks :).

OpenStudy (turingtest):

welcome, good luck :)

OpenStudy (konradzuse):

:D on to the next question!